OpenStudy (anonymous):

bridge built in shape of parabolic arch. Bridge arch span of 166 ft, max height 40ft. find height of arch @ 10ft from ctr. ok that website didn't help me that i was looking at. i understand it's a quadratic 2a=166 c=40 and i need to move x from ctr to start at 10 and i need that y value there. I must be getting tired

OpenStudy (anonymous):

ok figured it out. the equation is a83^2+40 (solving for a) a=-40/83^2 (half the distance of 166 is 83) f(10)=(-40/6889)(10^2)+40=39.4 ft must be tired. that was tricky to me. hope this helps someone if you are searching.